OpenStudy (anonymous):

What is the equation of the circle with center ___5, 2___ and radius ____4____? [Use different numbers] Explain each variable and number in the equation.

OpenStudy (lalaly):

lol why dont u do it @ParthKohli xD

Parth (parthkohli):

Okay! So, the equation of a circle is: \(\Large \color{MidnightBlue}{\Rightarrow (x - h)^2 + (y - b)^2 = r^2 }\) (h,b) are the center co-ordinates. r is the radius.

Parth (parthkohli):

Can you do it now @ScarletChains ?

OpenStudy (anonymous):

I hope so!! haha Thanks for the help!

OpenStudy (lalaly):

:D:D

OpenStudy (anonymous):

Okay, so @ParthKohli if h and b are center co-ordinates and r is the radius. x and y are.. just x and y? or do the stand for something else? (Sorry, guys, Geometry is NOT my thing!)

Parth (parthkohli):

Haha lana...I asked the same thing

OpenStudy (lalaly):

lol, if u want the equation and u have the center and the radius, u keep x and y as they are,,, but sometimes they give u (x-2)^2+(y-3)^2=r^2 and they tell u that it passes through a point (2,1) for example so find r ... to do that u just substitute (2,1) in x and y and then find r ... or they ask u to find the center,,, but in ur question,,, x and y are just x and y

OpenStudy (anonymous):

Thank you!!

OpenStudy (lalaly):

:):)

Parth (parthkohli):

Basically, x and y are the axis's

OpenStudy (lalaly):

yeah :D

OpenStudy (anonymous):

@ParthKohli and @lalaly so this is what I got.. I don't think I did it right!>< (x-h)^2 + (y-b)^2 = r^2 (x-5)^2 + (y-2)^2 = 4^2 (x-25) + (y-4) = 16 x + y – 29 = 16 x + y = 45

OpenStudy (lalaly):

noo just stop after ur second step the equation of the circle is (x-5)^2 + (y-2)^2 =16

OpenStudy (anonymous):

It's definitely official.. Geometry is not my calling! Thank you for all of your helpppp!